# The Rate-Distortion-Perception-Classification Tradeoff: Joint Source Coding and Modulation via Inverse-Domain GANs

### v1

@@ 1.0 $m=2$$m=8$$m=64$Refer to captionRate120120120120909090906060606030303030InputRefer to caption
Figure 9: (a) Rate-distortion curves of RDPCO for varying compressed dimension $m$; hyperparameters are $(P,C)=(4.1,0.1)$. (b) corresponding behavior of ID-GAN on MNIST images.
@@
Visual illustration on MNIST. Fig. 5 shows image examples from MNIST reconstructed by all the algorithms. In this experiment, the latent dimension $m$ was fixed to $8$ in D-JSCC [2], AE+GAN [8], and ID-GAN. The rate was then computed according to the expression in the objective of (7). For ID-GAN, we set $\lambda_{d}=\lambda_{p}=\lambda_{c}=1$. The images that the traditional system in unable to reconstruct the images even when the rate is high, likely due to the quantization and compression in JPEG. When the rate decreases, we observe a cliff effect, with the image quality degrading abruptly. D-JSCC [2], on the other hand, is based on an autoencoder which, given the high bandwidth ratio, outputs blurry images at all rates. Yet, despite the degraded perceptual quality, it still preserves semantic information. The images reconstructed by AE+GAN training [8], on the other hand, have poor perceptual quality for low rates ($1$ and $0.1$). This is likely because it is difficult to balance weights between multiple loss terms The proposed ID-GAN scheme, in contrast, not only preserves semantic information in images, but also reconstructs them with high perceptual quality. Compared to D-JSCC and AE+GAN, ID-GAN generates images with various styles, including different angles and thickness, even at extremely high bandwidth compression ratios.
V-B RDPCO framework
We now describe experiments using the RDPCO algorithm (Algorithm 2) to solve (37). Since the channel rate in JSCM depends on the latent dimension $m$ and noise levels in $\bm{\Sigma}$, we evaluate how these two factors influence the JSCM system. First, we fix $m$ and compute rate-distortion curves under different perception and classification error levels ($P$ and $C$); we visualize the effect on ID-GAN (Algorithm 1). Second, we investigate how the selection of $m$ influences the RDPC tradeoff for both RDPCO and ID-GAN.
@@ V-B RDPCO framework
Rate-distortion analysis. In this experiment, we vary both $P$ and $C$ in the constraints of (37). Fig. 6 shows rate-distortion curves for different pairs $(P,C)$. For a fixed $P$, decreasing $C$ increases the rate; similarly, for a fixed $C$, decreasing $P$ increases the rate as well. This validates the tradeoff established in Theorem 1. Fig. 6 shows, for different rates, a similar behavior of ID-GAN when we modify its hyperparameters accordingly. Note that the hyperparameters apply only during training; during testing, we inject different levels of channel noise, obtaining different rates. The first column weighs all the metrics equally. The second column imposes a stringer requirement on classification performance by setting $\lambda_{c}$ two orders of magnitude larger than $\lambda_{d}$ and $\lambda_{p}$ [akin to decreasing $C$ in (37)]. While classification accuracy is maintained, we observe loss of details, like digit orientation and line thickness. The third column imposes a larger weight on perception. At low rates, even though semantic information is altered, our algorithm still generates meaningful digits.
Metrics as a function of $\bm{D}$. Fig. 7 shows how distortion, perception, and classification error metrics vary with $D$ in (37). Here the input dimension is $n=5$ and the latent one is $m=2$. The metrics displayed in Fig. 7 are different from the ones in Fig. 4, namely the right-hand side of (27) for distortion, (29) for perception, and (31) for classification error. The reason we display these quantities is that RDPCO attempts to minimize them directly. In Figs. 77, we see that when $C$ (resp. $P$) is fixed, increasing $P$ (resp. $C$) increases either the distortion or perception metrics. This behavior is as expected according to the RDPC tradeoff. In Fig. 7, we observe that for a fixed $C$, modifying $P$ produces no significant effect on the classification error, indicating that the classification constraint becomes active before the perception one.
Effect of $\bm{m}$ on the rate-distortion tradeoff. Fig. 8 is similar to Fig. 7, but $P$ and $C$ are fixed, while the latent dimension $m$ varies. When $m=1$, all metrics are invariant to $D$. Indeed, in this case (cf. Fig. 8), the perception constraint is active and dominates the two other constraints. For $m=2,3$, their classification error behaves similarly, but $m=3$ achieves better perception and worse distortion.
Fig. 9 shows the rate-distortion curves under the same parameters as Fig. 8. We can see again that the rate is invariant to $D$ for $m=1$, since the perception constraint is the only active one. For $m=2,3$, the curves have the familiar tradeoff shape. In this case, $m=2$ yields a rate-distortion curve better than $m=3$. The reason is that, as we saw in Fig. 8, the perception constraint is the most stringent of three constraints, and $m=2$ leads to a better perception-distortion tradeoff. This translates into a better rate-distortion tradeoff in Fig. 9.
@@
Finally, Fig. 9 shows how $m$ and different rates affect the output images of ID-GAN. Images in the first column are blurry and discontinuous, indicating that they may disregard the perception and classification losses. However, when $m=8$, the output images seem to be more suitable for transmission at all rates, preserving perception, but with a few semantic mistakes, e.g., a digit $2$ becoming a $3$. When $m=64$, the algorithm requires higher rates to reconstruct the images accurately. These results corroborate the RDPC tradeoff we derived. Indeed, they point to the existence of an optimal latent dimension $m$ that minimizes the rate while satisfying the distortion, perception, and classification constraints.
VI Conclusions
@@
We formulated and analyzed the tradeoff between rate, distortion, perception, and rate (RDPC) in a joint source coding and modulation (JSCM) framework. We showed the existence of a tradeoff and proposed two algorithms to achieve it. One algorithm is based on inverse-domain GAN (ID-GAN) and works under a general scenario; the other algorithm is heuristic and was designed, under a series of simplifying assumptions, to minimize an upper bound on the RDPC function. Experimental results showed the superior performance of ID-GAN in comparison to a traditional method, in which source coding and modulation are designed separately, and to deep joint source-channel coding schemes. Experiments revealed that improving perception quality and classification accuracy require higher rates, and the existence of an optimal compressed/latent dimension that minimizes rate while satisfying constraints on distortion, perception, and classification.
@@ Appendix A Proof of Theorem 1
$\Big{(}p_{\bm{Y}|\bm{X}}^{(j)},\,p_{\widehat{\bm{X}}|\widehat{\bm{Y}}}^{(j)},\,\bm{\Sigma}^{(j)}\Big{)}$ :$\mathrel{\mathop{\ordinarycolon}}=\!\!\!\begin{array}[t]{cl}\underset{p_{\bm{Y}|\bm{X}},\,p_{\widehat{\bm{X}}|\widehat{\bm{Y}}},\,\bm{\Sigma}}{\text{argmin}}&\sum_{i=1}^{m}\log\big{(}1+\frac{1}{\bm{\Sigma}_{ii}}\big{)}\\ \text{s.t.}&\mathbb{E}\big{[}\Delta(\bm{X},\widehat{\bm{X}})\big{]}\leq D_{j}\\[2.84544pt] &d(p_{\bm{X}},\,p_{\widehat{\bm{X}}})\leq P_{j}\\[2.84544pt] &\epsilon_{c_{0}}(\bm{X},\widehat{\bm{X}})\leq C_{j}\,,\end{array}$ (51)
@@
$\geq\!\!\begin{array}[t]{cl}\underset{p_{\bm{Y}|\bm{X}},\,p_{\widehat{\bm{X}}|\widehat{\bm{Y}}},\,\bm{\Sigma}}{\min}&\sum_{i=1}^{m}\log\big{(}1+\frac{1}{\bm{\Sigma}_{ii}}\big{)}\\ \text{s.t.}&\mathbb{E}\big{[}\Delta(\bm{X},\widehat{\bm{X}})\big{]}\leq(1-\alpha)D_{1}+\alpha D_{2}\\ &d(p_{\bm{X}},\,p_{\widehat{\bm{X}}})\leq(1-\alpha)P_{1}+\alpha P_{2}\\ &\epsilon_{c_{0}}(\bm{X},\widehat{\bm{X}})\leq(1-\alpha)C_{1}+\alpha C_{2}\end{array}$ (57)
@@
$\mathcal{\epsilon}_{c_{0}}$ $(\bm{X},\,\widehat{\bm{X}}^{(\alpha)})=\sum_{i<j}p_{j}\cdot\int_{\mathcal{R}_{i}}\operatorname{d\!}\,p_{\widehat{\bm{X}}^{(\alpha)}|H_{j}}$ (66)
$=\sum_{i<j}p_{j}\cdot\int\int_{\mathcal{R}_{i}}\operatorname{d\!}\,p_{\widehat{\bm{X}}^{(\alpha)}|\widehat{\bm{Y}},\,H_{j}}\operatorname{d\!}\,p_{\widehat{\bm{Y}}}$ (67)
$=\sum_{i<j}p_{j}\cdot\int\int_{\mathcal{R}_{i}}\operatorname{d\!}\,p_{\widehat{\bm{X}}^{(\alpha)}|\widehat{\bm{Y}}}\operatorname{d\!}\,p_{\widehat{\bm{Y}}}$ (68)
$=(1-\alpha)\sum_{i<j}p_{j}\cdot\int\int_{\mathcal{R}_{i}}\operatorname{d\!}\,p_{\widehat{\bm{X}}^{(1)}|\widehat{\bm{Y}}}\operatorname{d\!}\,p_{\widehat{\bm{Y}}}$
$\qquad\qquad\qquad+\alpha\sum_{i<j}p_{j}\cdot\int\int_{\mathcal{R}_{i}}\operatorname{d\!}\,p_{\widehat{\bm{X}}^{(2)}|\widehat{\bm{Y}}}\operatorname{d\!}\,p_{\widehat{\bm{Y}}}$ (69)
$=(1-\alpha)\epsilon_{c_{0}}\big{(}\bm{X},\,\widehat{\bm{X}}^{(1)}\big{)}+\alpha\epsilon_{c_{0}}\big{(}\bm{X},\,\widehat{\bm{X}}^{(2)}\big{)}$ (70)

### v2

@@ VI-A RDPCO algorithm
Figure 4: Rate-distortion curves of RDPCO for (a) varying $P$ and $C$, and (b) varying compressed dimension $m$ with hyperparameters $(P,C)=(4.1,0.1)$.
Rate-distortion analysis. In this experiment, as in Fig. 2, we vary both $P$ and $C$ in the constraints of (32). Fig. 4 shows the resulting rate-distortion curves. For a fixed $P$, decreasing $C$ increases the rate; similarly, for a fixed $C$, decreasing $P$ increases the rate as well. This validates the tradeoff established in Theorem 1. Fig. 4 shows the rate-distortion curves under the same parameters as Fig. 3, i.e., $(P,C)=(4.1,0.1)$. We can see again that, for $m=1$, the rate is invariant to $D$, since the perception constraint is the only active one. For $m=2,3$, the curves have the familiar tradeoff shape. In this case, $m=2$ yields a rate-distortion curve better than $m=3$. The reason is that, as we saw in Fig. 3, the perception constraint is the most stringent of three constraints, and $m=2$ achieves a perception value closer to the limit of $4.1$, leading to a better rate-distortion tradeoff in Fig 4. These results corroborate the RDPC tradeoff we derived. Indeed, they point to the existence of an optimal latent dimension $m$ that minimizes the rate while satisfying the distortion, perception, and classification constraints.
[…]
kernel = 5
stride = 1
padding = 0
conv_transp1+ReLU
$\scriptstyle 7\times 7\times 128$
kernel = 5
stride = 1
padding = 0
conv_transp2
+ReLU
$\scriptstyle 14\times 14\times 64$
kernel = 5,  stride = 1
padding = 0
conv_transp3+ReLU
image
$\scriptstyle 28\times 28\times 1$
$\scriptstyle 14\times 14\times 64$
kernel = 5,  stride = 2
padding = 2
conv1+ReLU
$\scriptstyle 7\times 7\times 128$
kernel = 5
stride = 2
padding = 2
conv2+ReLU
$\scriptstyle 4\times 4\times 256$
kernel = 5
stride = 2
padding = 2
conv3+ReLU
$\scriptstyle 1\times 4096$FC1×111$\scriptstyle$ 1$\times$ 11 × 1(a)Decoder(b)Critics
[…]
0.96 Refer to caption28×28×128281$\scriptstyle$ 28$\times$ 28$\times$ 128 × 28 × 1imageFC1×7841784$\scriptstyle$ 1$\times$ 7841 × 784FC+BN+LeakyReLU1×5121512$\scriptstyle$ 1$\times$ 5121 × 5121×2561256$\scriptstyle$ 1$\times$ 2561 × 2561×1281128$\scriptstyle$ 1$\times$ 1281 × 128
FC+BN
+Tanh
[…]
Figure 6: Architecture of the encoder $e(\cdot\,;\,\bm{\theta}_{e})$, consisting of fully connected layers of indicated dimensions. Layers 2-4 contain a batch normalization (BN) layer and use LeakyReLU as activation. Layer 5 uses $\tanh$ instead.
Refer to captionMean-squared error (MSE)Rate-10 dB5 dB20 dB
Refer to captionFréchet inception distance (FID)Rate-10 dB5 dB20 dB
Refer to captionClassification errorRate-10 dB5 dB20 dB
Figure 7: Different metrics versus rate (41) for the proposed ID-GAN, D-JSCC [deep-jscc], and AE+GAN [agustsson2019generative]. In ID-GAN, we set $(\lambda_{g},\lambda_{d},\lambda_{p},\lambda_{c})=(1,10^{3},1,10^{3})$ and, during training, we either learned $\text{SNR}_{t}:=-10\log_{10}\sigma_{t}^{2}$ or fixed it to $-10$ dB, $5$ dB, $20$ dB (vertical lines depicting the corresponding rates), or $\infty$ (no noise). For all the metrics, the lower the better. (a) mean-squared error (MSE), (b) Fréchet inception distance (FID), and (c) classification error.
[…]
Figure 8: Example images reconstructed by a traditional method (JPEG+LDPC+BPSK), D-JSCC, AE+GAN training, and proposed ID-GAN with learned $\text{SNR}_{t}$ and parameters $(\lambda_{g},\lambda_{d},\lambda_{p},\lambda_{c})=(1,500,1,10^{3})$.
[…]
@@ 1.0 $m=2$$m=8$$m=64$Refer to captionRate120120120120909090906060606030303030InputRefer to caption
Figure 9: (a) Corresponding behavior of ID-GAN on MNIST images. The triples on top of each column represent hyperparameters $(\lambda_{d},\lambda_{p},\lambda_{c})$. We fixed $\text{SNR}_{t}=\infty$. (b) corresponding behavior of ID-GAN on MNIST images with fixed $\text{SNR}_{t}=\infty$.
Refer to captionInputRefer to captionMSE loss(ID-GAN)Refer to captionMSE + CE loss(ID-GAN)Refer to captionID-GAN
Figure 10: Visualization of the results of the ablation study. In the second column, only the MSE term of (47) was considered during training. In the third, both the MSE and the cross-entropy terms were considered. And in the fourth, the full loss (47) was considered.
VI-B ID-GAN algorithm
We now consider ID-GAN (Algorithm 2) applied to the popular MNIST dataset [LeCun98], which has $60$k training images and $10$k test images of size $28\times 28$ depicting digits from 00 to $9$ ($10$ classes). Before describing the experimental setup in detail, we explain how all the functions in the ID-GAN framework in Fig. 1 were implemented.
Network architectures. Fig. 5 shows the architectures of the decoder $d(\cdot\,;\,\bm{\theta}_{d})$ [Fig. 5(a)] and of the critics $f_{1}$ and $f_{2}$ [Fig. 5(b)]. The latter have the same architecture, but they are initialized independently, with different seeds. The decoder in Fig. 5(a) increases the dimensions of the data by first using a fully connected network, whose output is reshaped to a $4\times 4\times 256$ tensor, and then by upsampling along the channel. The upsampling is performed via transposed convolutions with ReLU activations. The architecture of the critics [Fig. 5(b)] is symmetric to that of the decoder. The input image is compressed via a convolutional network, whose output in the last layer is mapped to a probability vector with a sigmoid function. The architecture of the encoder, on the other hand, downsamples the input image using only fully connected layers, as shown in Fig. 6. The network architecture of the classifier [cf. Fig. 1, bottom] is similar to the one of the encoder, except that the last layer is mapped to a 10-dimensional vector (coinciding with the number of classes of MNIST) and a sigmoid is applied to each entry. We train the classifier beforehand and fix it when training the encoder.
Experimental setup. In Algorithm 2, we used a learning rate $\alpha=10^{-5}$ and acceleration parameters $\beta_{1}=0.5$ and $\beta_{2}=0.9$ for Adam, a batch size of $S=50$, and $n_{\text{critic}}=5$ iterations for the inner loop of the critic. The loss hyperparameters and training noise level $\text{SNR}_{t}$ will be reported for each experiment. Also, we ran the algorithm for just $5$ epochs. The reason for such a small number is that, as described in Section V-C, the decoder has already been trained when we run Algorithm 2. In fact, the decoder was also trained with few epochs, $8$.
Algorithms. We compare the proposed ID-GAN algorithm against D-JSCC [deep-jscc], the parallel autoencoder-GAN (AE+GAN) [agustsson2019generative], and a traditional approach in which source coding, channel coding, and modulation are designed separately: JPEG and Huffman codes for source coding, $3/4$ LDPC codes for channel coding, and BPSK for modulation [Gallager62-LDPC].
Metrics and comparison. For comparison metrics, we selected the mean squared error (MSE), the Fréchet inception distance (FID) [heusel2017gans], and the classification error $(1/|\mathcal{V}|)\sum_{v\in\mathcal{V}}\mathbbm{1}_{c_{0}(\widehat{\bm{x}}^{(v)})\neq\bm{\ell}^{(v)}}$, where $\mathcal{V}$ is the test set, and $\mathbbm{1}$ the 00-$1$ loss. FID captures perception quality by measuring the similarity between distributions of real and generated images. The smaller the FID, the closer the distributions. In summary, the smaller all the metrics, the better the performance. Comparing the performance of a JSCM system against a traditional system, however, is challenging. For example, [deep-jscc] proposed to use the ratio of bandwidth compression. Yet, in a traditional system, it is not obvious how to accurately determine the ratio between the size of images and the corresponding vectors in IQ-domain. Instead, we will use rate as defined in (41), i.e., channel rate, which quantifies the amount of information that a symbol can convey through a channel. In particular, the dimension $m$ of the latent variable $\bm{y}$ corresponds to the number of constellations in a traditional system. Thus, to compare the different algorithms in a fair way, we fix during testing the channel rate (41), making the results independent from the latent dimension $m$.
Results. Fig. 7 shows how the above metrics vary with the rate (41), measured in bits/image, for D-JSCC [deep-jscc], AE+GAN [agustsson2019generative], and the proposed ID-GAN. The curves represent the average values over the $10$k test images of MNIST. For ID-GAN, we fixed the hyperparameters as $(\lambda_{g},\lambda_{d},\lambda_{p},\lambda_{c})=(1,10^{3},1,10^{3})$ and considered four different fixed training noise levels, $\text{SNR}_{t}\in\{-10,5,20,\infty\}$ dB [$\text{SNR}_{t}=\infty$ corresponds to no noise], and also a value of $\text{SNR}_{t}$ that is learned during training. The vertical lines in Figs. 7-7 indicate the rates corresponding to $\text{SNR}_{t}=-10$ dB, $5$ dB and $20$ dB.
Fig. 7 indicates that image distortion, measured by the MSE, decreases with the rate for all the algorithms. D-JSCC outperforms all the other methods, as it was designed to minimize image distortion (MSE). Indeed, according to the RDPC tradeoff (Theorem 1), if the perception and classification metrics are ignored, an algorithm can achieve a smaller distortion for a given rate. When ID-GAN is trained with fixed $\text{SNR}_{t}$, we observe that the larger the value of $\text{SNR}_{t}$, the lower the achieved MSE for all rates. As we need to take into account other metrics (FID, classification error), this indicates the difficulty of manually selecting $\text{SNR}_{t}$. The version of ID-GAN with learned $\text{SNR}_{t}$ found an optimal value of $\text{SNR}_{t}=16.5$ dB during training, and its MSE performance follows closely the performance of ID-GAN versions with $\text{SNR}_{t}=20$ dB and $\text{SNR}_{t}=\infty$ (the lines of latter practically coincide).
In Fig. 7, however, D-JSCC is the algorithm with the worst FID performance, with an optimal point around a rate of 5 bits/image. For AE+GAN, the FID metric decreases monotonically, eventually surpassing ID-GAN trained with $\text{SNR}_{t}=20$ dB and $\infty$. Indeed, these versions of ID-GAN were the best overall in terms of FID performance, again closely followed by ID-GAN with learned $\text{SNR}_{t}$. When ID-GAN is trained with $\text{SNR}_{t}=-10$ dB or $5$ dB, FID reaches a minimum around a rate of 10 bits/image and then it increases. When $\text{SNR}_{t}$ is learned during training, its FID value decreases until a rate of 20 bits/image, and then it increases gradually. This implies that when $\text{SNR}_{t}$ is learned, the perception quality of ID-GAN is stable for a certain range of the rate. The reason why FID values increase for larger rates in ID-GAN is likely due to the mismatch between the noise levels during testing and training: indeed, the smaller the training noise level, the earlier the FID curve starts to increase. This happens only for FID likely because the decoder, which is largely responsible for the perception quality of the output, is fixed during the training of the encoder. This decoupling may make perception quality more sensitive to discrepancies between training/testing setups.
Fig. 7 shows how the classification error of the algorithms varies with the rate. All the versions of ID-GAN [except for $\text{SNR}_{t}=-10$ dB, for rate $>40$ bits/image] outperform both D-JSCC and AE+GAN, which have almost overlapping lines. Within ID-GAN versions with fixed $\text{SNR}_{t}$, we see that $\text{SNR}_{t}=5$ dB yields the best classification error across all the rates. Decreasing it to $-10$ dB or increasing it to $20$ dB or $\infty$ results in a worse classification error. The downside of such a good performance in terms of classification error is the poor performance in terms of distortion [Fig. 7] and perception [Fig. 7] for large rates. ID-GAN with learned $\text{SNR}_{t}$, on the other hand, achieves a good performance across all metrics [Figs. 7-7], effectively balancing rate and all the metrics. We conclude that ID-GAN achieves a tradeoff better than AE+GAN, outperforming it in all metrics for rates $\lesssim 40$ bits/image. It also achieves perception and classification performance better than D-JSCC, at the cost of worse distortion.
@@ VI-B ID-GAN algorithm
Visual illustration. Fig. 8 shows concrete image examples reconstructed by all the algorithms. In this experiment, the latent dimension $m$ in D-JSCC [deep-jscc], AE+GAN [agustsson2019generative], and ID-GAN was fixed to $8$ and the respective rate was then computed via (41). The parameters of ID-GAN were set to $(\lambda_{g},\lambda_{d},\lambda_{p},\lambda_{c})=(1,500,1,10^{3})$, and $\text{SNR}_{t}$ was learned. The first column depicts the images reconstructed by a traditional JPEG+LDPC+BPSK system. Even when the rate is high, the reconstructed images fail to match the input one, likely due to quantization and compression artifacts in JPEG. As the rate decreases, we observe a cliff effect, with the image quality degrading abruptly. D-JSCC [deep-jscc] in the second column, on the other hand, is based on an autoencoder which, given the high bandwidth ratio, outputs blurry images at all rates. As in Fig. 7, even though D-JSCC outperforms the other algorithms in MSE, it has a relatively poor perception quality (FID) and classification error. In the third column, the images reconstructed by AE+GAN [agustsson2019generative] have poor perceptual quality for a low rate, i.e., $10$. We noticed that the quality of its output images collapses when its training parameters $\lambda_{d}$, $\lambda_{p}$, and $\lambda_{c}$ are not carefully set, revealing the difficulty in balancing different loss terms. The proposed ID-GAN scheme shown in the fourth column, in contrast, not only preserves semantic information in images, but also reconstructs them with high perceptual quality. Compared to D-JSCC and AE+GAN, ID-GAN generates images with various styles, including different angles and thickness, even at extremely high bandwidth compression ratios.
@@
Rate-distortion analysis. To illustrate how ID-GAN and RDPCO behave similarly, Fig. 9 adapts to MNIST the setup of Fig. 4, which shows rate-distortion curves of RDPCO with varying $(P,C)$ or $m$. We fixed the training noise level of ID-GAN to $\text{SNR}_{t}=\infty$ so that we can visualize the effect of modifying its hyperparameters. Fig. 9 shows, for different rates, a behavior similar to RDPCO in Fig. 4 when we modify its hyperparameters $(\lambda_{d},\lambda_{p},\lambda_{c})$ accordingly. Note that these hyperparameters apply only during training; during testing, we inject different levels of channel noise, obtaining different rates. The first column of Fig. 9 weighs all the metrics equally. The second column imposes a more stringent requirement on classification performance by setting $\lambda_{c}$ two orders of magnitude larger than $\lambda_{d}$ and $\lambda_{p}$ [akin to decreasing $C$ in (32)]. While classification accuracy is maintained, we observe loss of details, like digit orientation and line thickness. The third column imposes a larger weight on perception. At low rates, even though semantic information is altered, our algorithm still generates meaningful digits.
@@
Fig. 9, akin to Fig. 4, shows how different values of $m$ affect the output images of ID-GAN for different rates. Images in the first column, for $m=2$, are blurry and discontinuous, indicating that they may disregard the perception and classification losses. However, when $m=8$, the output images seem to be more suitable for transmission at all rates, preserving perception, but with a few semantic mistakes, e.g., a digit $2$ becoming a $3$. When $m=64$, the algorithm requires higher rates to reconstruct the images accurately.
Ablation study. We performed a small ablation study to assess whether all the terms in the loss associated to the encoder (47) are necessary for good performance. To do so, we fixed the noise level at $\text{SNR}_{t}=30$ dB, corresponding to a rate of $50$, which is large enough to enable seeing visual differences in the reconstructed images. First, we trained the encoder with just the MSE loss, i.e., $(\lambda_{d},\lambda_{p},\lambda_{c})=(1,0,0)$ in (47). A set of reconstructed digits from the test set is shown in the second column of Fig. 10. The digits are blurry, as the encoder learns just to compress pixel-level information and disregards any semantic information. Then, we trained the encoder with the MSE and classification loss, i.e., $(\lambda_{d},\lambda_{p},\lambda_{c})=(1,0,40)$. As shown in the third column of Fig. 10, the algorithm preserved semantic information better, e.g., correctly depicting a $4$ in the $3$rd row and column, but the digits exhibit poor diversity. For example, all the 2’s look similar. Finally, we trained the encoder using the full loss, with $(\lambda_{d},\lambda_{p},\lambda_{c})=(1,1,40)$. The reconstructed digits in the last column of Fig. 10 are not only sharp, of the correct class (except for the $4$ in row 1, column 2) but also exhibit more diversity, thus looking more realistic.
VII Conclusions
@@
We formulated and analyzed the tradeoff between rate, distortion, perception, and rate (RDPC) in a joint source coding and modulation (JSCM) framework. We showed the existence of a tradeoff and proposed two algorithms to achieve it. One algorithm is heuristic and was designed under simplifying assumptions to minimize an upper bound on the RDPC function; the other was based on inverse-domain GAN (ID-GAN) and works under a general scenario. Experimental results showed that ID-GAN achieves a better tradeoff than a traditional method, in which source coding and modulation are designed separately, and a tradeoff better or similar to recent deep joint source-channel coding schemes. Experiments revealed that improving perception quality and classification accuracy require higher rates, and also showed the existence of an optimal compressed/latent dimension that minimizes rate while satisfying constraints on distortion, perception, and classification.
Acknowledgments
We thank the two reviewers for their insightful suggestions, which significantly improved the quality of the paper.
[…]
@@ Appendix A Proof of Theorem 1
$\Big{(}p_{\bm{Y}|\bm{X}}^{(j)},\,p_{\widehat{\bm{X}}|\widehat{\bm{Y}}}^{(j)},\,\bm{\Sigma}^{(j)}\Big{)}$ $:=\!\!\!\begin{array}[t]{cl}\underset{p_{\bm{Y}|\bm{X}},\,p_{\widehat{\bm{X}}|\widehat{\bm{Y}}},\,\bm{\Sigma}}{\text{argmin}}&\sum_{i=1}^{m}\log\big{(}1+\frac{1}{\bm{\Sigma}_{ii}}\big{)}\\ \text{s.t.}&\mathbb{E}\big{[}\Delta(\bm{X},\widehat{\bm{X}})\big{]}\leq D_{j}\\[2.84544pt] &d(p_{\bm{X}},\,p_{\widehat{\bm{X}}})\leq P_{j}\\[2.84544pt] &\mathbb{E}\big{[}\epsilon_{c_{0}}(\bm{X},\widehat{\bm{X}})\big{]}\leq C_{j},\end{array}$ (54)
@@
$\geq\!\!\begin{array}[t]{cl}\underset{p_{\bm{Y}|\bm{X}},\,p_{\widehat{\bm{X}}|\widehat{\bm{Y}}},\,\bm{\Sigma}}{\min}&\sum_{i=1}^{m}\log\big{(}1+\frac{1}{\bm{\Sigma}_{ii}}\big{)}\\ \text{s.t.}&\mathbb{E}\big{[}\Delta(\bm{X},\widehat{\bm{X}})\big{]}\leq(1-\alpha)D_{1}+\alpha D_{2}\\ &d(p_{\bm{X}},\,p_{\widehat{\bm{X}}})\leq(1-\alpha)P_{1}+\alpha P_{2}\\ &\mathbb{E}\big{[}\epsilon_{c_{0}}(\bm{X},\widehat{\bm{X}})\big{]}\leq(1-\alpha)C_{1}+\alpha C_{2}\end{array}$ (60)
@@
[…]
$=(1-\alpha)\mathbb{E}\big{[}\epsilon_{c_{0}}\big{(}\bm{X},\,\widehat{\bm{X}}^{(1)}\big{)}\big{]}+\alpha\mathbb{E}\big{[}\epsilon_{c_{0}}\big{(}\bm{X},\,\widehat{\bm{X}}^{(2)}\big{)}\big{]}$ (73)